A stock is currently selling for $68.96. A call option expiring in 194 days with a strike price of $65.00 is selling for $8.12; A put option expiring in 133 days with a strike price of $70.00 is selling for $4.57. What is the call option's exercise value? a. $0.00 b. $0.20 c. $0.43 d. $0.61 e. $0.69 f. $1.04 g. $2.22 h. $2.49 i. $3.12 j. $3.53 k. $3.55 l. 3.96 m. $4.16 n. $4.57 o. $5.00 p. $6.74 q. $7.08 r. $7.49 s. $7.69 t. $8.12 u. $12.69 v. none of the above

A stock is currently selling for $68.96. A call option expiring in 194 days with a strike price of $65.00 is selling for $8.12; A put option expiring in 133 days with a strike price of $70.00 is selling for $4.57. What is the call option's time premium? a. $0.00 b. $0.20 c. $0.43 d. $0.61 e. $0.69 f. $1.04 g. $2.22 h. $2.49 i. $3.12 j. $3.53 k. $3.55 l. $3.96 m. $4.16 n. $4.57 o. $5.00 p. $6.74 q. $7.08 r. $7.49 s. $7.69 t. $8.12 u. $12.69 v. none of the above

A stock is currently selling for $68.96. A call option expiring in 194 days with a strike price of $65.00 is selling for $8.12; A put option expiring in 133 days with a strike price of $70.00 is selling for $4.57. What is the put option's exercise value? a. $0.00 b. $0.20 c. $0.43 d. $0.61 e. $0.69 f. $1.04 g. $2.22 h. $2.49 i. $3.12 j. $3.53 k. $3.55 l. $3.96 m. $4.16 n. $4.57 o. $5.00 p. $6.74 q. $7.08 r. $7.49 s. $7.69 t. $8.12 u. $12.69 v. none of the above
Answer #1

Exercise value of call=MAX(spot-strike,0)

Exercise value of put=MAX(strike-spot,0)

Time value=Premium-Exercise value

1.

=MAX(68.96-65,0)=3.96

2.

=8.12-MAX(68.96-65,0=4.16

3.

=MAX(70-68.96,0=1.04